# SeSE: A Structural Information-Guided Uncertainty Quantification Framework for Hallucination Detection in LLMs

### v2

@@ VIII-F1 Response sampling prompt
VIII-F2 Claim decomposition prompt
Given a LLM response, the claim decomposition prompt is used to break down the generated response into a set of individual claims. The prompt is as follows:
    The input is:
    {greedily decoded response}
    Output:
VIII-F3 Ground-truth evaluation prompt
For sentence-length generation, we automatically determine whether the given answer is correct or incorrect by comparing it with the reference answer using GPT-5-mini. We use the following template:
@@ VIII-F3 Ground-truth evaluation prompt
For the dataset SQuAD_V2 with multiple reference answers, the second line becomes “The following are the reference answers:”, and the last line asks “determine whether the proposed answer has the same meaning as any of the reference answers within the context of the question.”
VIII-G Dataset and Annotation Details
@@
SeSE can detect confabulations in free-form text generation across a range of domains without requiring prior domain knowledge. We evaluate it on sentence-level question-answering tasks spanning life sciences, mathematical reasoning, trivia knowledge, open-domain QA, and commonsense reasoning. Furthermore, to examine SeSE’s hallucination detection capability in long-text generation tasks, we construct four custom datasets based on the generated outputs of DeepSeek‑V3.1 and Gemini‑2.5‑Flash on two benchmark datasets—FActScore and PopQA. Since current LLMs have an excessively high accuracy rate when using tools, all experiments are conducted in offline mode, relying solely on their own capabilities. In Table. VI, we show samples and statistics for each dataset. All experimental datasets have been made publicly available in the code repository to facilitate reproducibility and further development.
VIII-G1 Datasets in sentence-length experiments
@@
In the sentence-length experiment, we utilize five representative QA datasets covering different domains: BioASQ [37], SVAMP [38], TriviaQA [39], NQ-Open [40], and SQuAD_V2 [41]. BioASQ derives from the annual biomedical semantic-indexing and question-answering challenge of the same name, focusing on life sciences. We select dataset from Task B of the 2023 BioASQ challenge. SQuAD_V2, a reading comprehension dataset, contains answers extracted from Wikipedia paragraphs. We exclude unanswerable questions designed to induce erroneous responses. TriviaQA encompasses trivia questions across multiple domains, including history, science, entertainment and so on. SVAMP features elementary mathematical word problems that test reasoning abilities. NQ-Open, an open-domain subset of Natural Questions, comprises real Google search queries with answers found in Wikipedia documents.
For each dataset, we sample 300 examples for training and 300 for testing each time with different random seed. Notably, SeSE is completely independent of the training data. For datasets with predefined partitions, sampling occurred within their respective training and test sets. All experiments employ free-form question answering rather than multiple-choice or true/false formats to better assess the characteristics of LLMs in generating free-from text. As the questions become too easy for the current LLMs when context is provided, we withhold context paragraphs across all experiment datasets except SVAMP to increase difficulty and provoke potential hallucinations.
TABLE VI: Details of sentence-length experiment datasets.
Dataset (Linked) Domain Count Example question Context Answer
BioASQ Life sciences 4.7k What disease is the ALK tyrosine kinase associated with? *Discard Cancer.
SVAMP Mathematical word problems 1k How many white t-shirts will she have? If mom buys 57 packages of white t-shirts and 34 trousers. White t-shirts can be purchased in packages of 53. 3021.
TriviaQA Trivia knowledge 15.4k Claude Littner is the new face in the latest series of which programme? *Discard Apprentice.
NQ-Open Open-domain natural questions 91.5k The removal of temperature in fire fighting method is known as? *Discard Cooling.
SQuAD General knowledge 141.9k When was Otto von Bismarck born? *Discard 1862.
FactScore Biographies 3.5k Tell me a bio of Maxime Masson.
PopQA Multi-topic Wikipedia entries 3.8k Provide me with a paragraph detailing some facts related to Speak (2004 film).
  • *Discard denotes that we don’t include context in prompts.
VIII-G2 Datasets in long-form experiments
@@
We select DeepSeek-V3.1 and Gemini-2.5-Flash as representative SOTA LLMs at the time of writing. Due to the absence of fine-grained, long-form hallucination evaluation datasets for these two models, we construct four new datasets based on entities from FActScore [27] and PopQA [42].
@@
FActScore: This dataset is widely used for evaluating the factuality of biographies generated by LLMs, with entities sourced from Wikipedia. We select entities from the publicly released “unlabelled” portion and prompt each model with the query “Tell me a bio of [subject]” to collect long-form generations.
@@
Annotation Process: For each model-dataset pair (four pairs in total), we employ a two-stage cross-validation annotation process. First, we generate factual claims for each sampled entity, following the processing pipeline detailed in Section III-C. Each claim is then evaluated for factual accuracy according to the following criteria:
  • True: The claim is factually accurate and can be verified by the corresponding Wikipedia page.
@@
To enhance annotation efficiency and consistency, we utilize two powerful reasoning models, GPT-4 and DeepSeek-R1, for automated labeling. A claim is automatically assigned a label only if both models produce identical classifications; otherwise, the claim is reviewed by human annotators for final decision. This hybrid annotation framework ensures high accuracy while enabling large-scale processing. In total, 7,407 generated claims have been annotated through this process.
We find that even such powerful LLMs exhibit significant hallucination rates, with 28% for DeepSeek-V3.1 and 27% for Gemini-2.5-Flash. Although hallucinations, like software vulnerabilities, cannot be completely eliminated at the current stage, effective detection mechanisms can mitigate associated risks and enhance the reliability of LLMs. Given that hallucinations may lead to serious real-world consequences, reliable uncertainty quantification is crucial for ensuring the responsible deployment of LLM-based systems, which highlights the important value of our research.
VIII-H Baselines Details
@@
1) Baselines in sentence-length experiments
Embedding Regression (ER) [24]. Embedding Regression represents the typical supervised learning approach. Kadavath et al. [24] develop predictors by fine-tuning proprietary language models on annotated QA datasets to assess whether target LLMs could correctly answer specific questions. Our experiment implements a more efficient alternative: we directly extract the final hidden layer states from LLMs and train a Embedding Regression classifier to achieve equivalent predictive functionality without requiring model fine-tuning or ground-truth answers.
P(True) [24]. This method first prompts LLMs to generate multiple distinct answers, then presents this answer list alongside the greedy decoding response and a binary question: “Is this answer (a) correct or (b) incorrect?”. The confidence score is derived from the probability with which the model selects option “(a)”. We enhance P(True) through few-shot prompting by incorporating ten randomly selected training examples formatted according to the described protocol and annotated with their true labels. This strategy represents a form of supervised in-context learning that leverages partial ground truth without necessitating model retraining.
@@
Length-normalized Predictive Entropy (LN-PE) [47]. To calculate prediction entropy, we must obtain the probabilities that LLMs assign to generated token sequences. The probability of an entire sequence $s$ given context $x$ equals the product of probabilities for each token conditioned on previous tokens, with its log probability expressed as $\log P(s|x)=\sum_{i}\log P(s_{i}|s_{<i},x)$. $s_{i}$ is the $i$-th output token and $s_{<i}$ denotes all preceding tokens. Due to the conditional independence of token probabilities, longer sequences inherently have lower joint likelihood. The joint likelihood of a sequence decreases exponentially with length $L$, while its negative log probability increases linearly with $L$, causing longer sentences to contribute disproportionately to entropy. Length-normalized prediction entropy addresses this bias by normalizing the log probability by sequence length, using the arithmetic mean $\frac{1}{L}\sum_{i=1}^{L}\log P(s_{i}|s_{<i},x)$. This normalization effectively assumes that the uncertainty of the generated result is independent of sequence length.
@@
Semantic Entropy (SE) and Discrete Semantic Entropy (DSE) [14, 16]. Semantic entropy aims to evaluate the uncertainty of LLMs regarding the meaning of their generated sequences. It first calculates the sum of probabilities for all token sequences that can be considered as expressing the same meaning. Given context $x$, for each semantic equivalence class $c$, its probability $P(c|x)$ is estimated through semantic clustering of generated sequences $s$:
@@
where $C_{i}$ are $M$ clusters extracted from the NN generations and $p^{\prime}(C_{i}\mid x)=\frac{p(C_{i}\mid x)}{\sum_{k}p(C_{k}\mid x)}$, which we refer to as $p(C_{i}\mid x)$ in the following for simplicity. DSE can be extended to cases where token likelihoods are not available by approximating $p(C_{i}\mid x)$ with the fraction of generated texts in each cluster, $p(C_{i}\mid x)\approx\sum_{i=1}^{N}\mathbb{I}(s_{i}\in C_{i})/N$.
Kernel Language Entropy (KLE)[17]. To address the limitations of traditional semantic entropy, which only rely on semantic equivalence relations and fail to capture fine-grained semantic similarities, Nikitin et al. proposed Kernel Language Entropy (KLE). Its encodes the semantic similarities between texts generated by LLMs through a unit trace positive semidefinite kernel $K_{\text{sem}}$, and then use the von Neumann Entropy (VNE) to quantify the uncertainty of the semantic space represented by this kernel. For a unit trace positive definite matrix $A\in\mathbb{R}^{n\times n}$, its von Neumann Entropy, denoted as $\text{VNE}(A)$, is defined in the form of matrix trace operation. On this basis, the KLE is defined as the VNE of the semantic kernel:
@@
SelfCheck-Prompt (SC-Prompt) [25]. SC-Prompt represents the highest-performing variant within the SelfCheckGPT framework. Its central idea is to prompt the LLM to assess the semantic consistency between the target sentence and a set of randomly generated samples, thereby determining the presence of hallucinations. The procedure is as follows: given a sentence to be evaluated $r_{i}$ and $N$ randomly generated samples $S^{n}$ corresponding to the same query, a fixed prompt template is used to instruct the LLM to evaluate their semantic consistency. The LLM’s output is then converted into a numerical score $x$ in accordance with predefined rules. Finally, the inconsistency score for the sentence is obtained by averaging these individual scores. A score approaching 1.0 indicates a higher likelihood that the sentence contains hallucinations. The calculation formula is:
${SC}_{Prompt}=\frac{1}{N}\sum_{n=1}^{N}x^{n},$ (28)
where $x_{i}^{n}$ is the mapping score corresponding to the $n$-th sample.
2)Baselines in long-form experiments
Discrete Semantic Entropy (DSE) variant [16]. For long-form generation, discrete semantic entropy operates through three key steps: (1) decomposing the text into atomic claims, (2) generating multiple possible questions that could trigger each claim in reverse, and (3) querying the original LLMs to produce new responses for each question, while including the original claim as a candidate. The final uncertainty estimate for each claim is derived by averaging the semantic entropy values across all associated questions. In our long-form experiments, we implement discrete semantic entropy according to the original paper’s best practices. Applying discrete semantic entropy to long-form generation introduces additional assumptions and complexities, and its computational cost increases with higher sampling. In contrast, SeSE achieves better performance with a more concise principle and lower cost.
Verbalized confidence (VC) [22]. Verbalized Confidence prompts LLMs to directly express their confidence in a claim through natural language. Uncertainty is quantified by parsing the confidence expressions in the LLMs’ output (e.g., “very confident”, “100%”, etc.) and mapping them to numerical values. Here, we mainly consider two variants:
Post-hoc Verbalized Confidence (PH-VC): This method elicits the verbalized confidence in a post-hoc manner after the entire claim set $C$ has been decomposed from generations. Specifically, we prompt an LLM to express its confidence about each claim $c\in C$ given multiple options such as “Unlikely (40%)”, “Even chance (50%)” etc. The specific prompt that we use is as following:

### v3

@@ E.1 Response sampling prompt
Table 5 includes two examples showing the sampled responses of the Llama-3.1-8B to two questions.
Table 5: Example Responses from BioASQ and NQ-Open datasets sampled from Llama-3.1-8B.
Example Responses from BioASQ Example Responses from NQ-Open
Input: Answer the following question concisely in one sentence. Question: Which organs are mostly affected in Systemic Lupus Erythematosus (SLE)? Answer: Input: Answer the following question concisely in one sentence. Question: Which financial statement involves all aspects of the accounting equation? Answer:
Reference Answer Reference Answer
kidney, brain, heart, skin balance sheet
Greedily-Decoded Response Greedily-Decoded Response
SLE primarily affects the skin, kidneys, and heart. accounting equation.
Randomly Sampled Responses Randomly Sampled Responses
$r^{(1)}$: including the skin, joints, kidneys, and kidneys $r^{(2)}$: the skin, joints, kidneys, and heart. $r^{(3)}$: the skin, joints, kidneys, heart, brain, and kidneys. $r^{(4)}$: the skin, joints, kidneys, and the immune system. $r^{(5)}$: the skin, joints, kidneys, and lymph nodes. $r^{(6)}$: the joints, skin, kidneys, and kidneys. $r^{(7)}$: the skin, kidneys, heart, brain, and immune system. $r^{(8)}$: the skin, kidneys, and lupus-prone joints. $r^{(9)}$: the skin, joints, kidneys, heart, and brain. $r^{(10)}$: the skin, kidneys, heart, and lungs. $r^{(1)}$: Balance Sheet. $r^{(2)}$: the balance sheet. $r^{(3)}$: balance sheet. $r^{(4)}$: balance sheet. $r^{(5)}$: Income statement. $r^{(6)}$: the statement of cash. $r^{(7)}$: the cash flow statement. $r^{(8)}$: the Balance sheet. $r^{(9)}$: the statement of equity or stockholders’ equity. $r^{(10)}$: the cash flow statement.
E.2 Ground-truth evaluation prompt
For short-form generation, we automatically determine whether the given answer is correct or incorrect by comparing it with the reference answer using GPT-5-mini. The specific prompt that we adapted from Farquhar et al. [2024] is as following:
@@ E.2 Ground-truth evaluation prompt
For the dataset SQuAD_V2 with multiple reference answers, the second line becomes “The following are the reference answers:", and the last line asks “determine whether the proposed answer has the same meaning as any of the reference answers within the context of the question."
Appendix F Details of Used Datasets
@@
SeSE can detect confabulations in free-form text generation across a range of domains without requiring prior domain knowledge. We evaluate it on short-form question-answering tasks spanning life sciences, mathematical reasoning, trivia knowledge, open-domain QA, and commonsense reasoning. Furthermore, to examine SeSE’s hallucination detection capability in long-text generation tasks, we construct four custom datasets based on the generated outputs of DeepSeek-V3.1 and Gemini-3-Flash on two benchmark datasets—FActScore and PopQA. Since current LLMs have an excessively high accuracy rate when using tools, all experiments are conducted in offline mode, relying solely on their own capabilities. All experimental datasets have been made publicly available in the code repository to facilitate reproducibility and further development.
F.1 Datasets in short-form experiments
@@
In the short-form experiment, we utilize five representative QA datasets covering different domains: BioASQ Krithara et al. [2023], SVAMP Patel et al. [2021], TriviaQA Joshi et al. [2017], NQ-Open Kwiatkowski et al. [2019], and SQuAD_V2 Rajpurkar [2018]. BioASQ derives from the annual biomedical semantic-indexing and question-answering challenge of the same name, focusing on life sciences. We select dataset from Task B of the 2023 BioASQ challenge. SQuAD_V2, a reading comprehension dataset, contains answers extracted from Wikipedia paragraphs. We exclude unanswerable questions designed to induce erroneous responses. TriviaQA encompasses trivia questions across multiple domains, including history, science, entertainment and so on. SVAMP features elementary mathematical word problems that test reasoning abilities. NQ-Open, an open-domain subset of Natural Questions, comprises real Google search queries with answers found in Wikipedia documents. For each dataset, we sample 300 examples for training and 300 for testing each time with different random seed. Notably, SeSE is completely independent of the training data. For datasets with predefined partitions, sampling occurred within their respective training and test sets. All experiments employ free-form question answering rather than multiple-choice or true/false formats to better assess the characteristics of LLMs in generating free-from text. As the questions become too easy for the current LLMs when context is provided, we withhold context paragraphs across all experiment datasets except SVAMP to increase difficulty.
Below are some example prompts of the short-form datasets we used, which are feed to the LLMs according to the template described in Appendix E.
BioASQ Answer the following question concisely in one sentence.
Question: What is the msDNA?
Answer:
Reference Answer:
msDNA is actually a complex of DNA, RNA, and probably protein.
NQ-Open Answer the following question concisely in one sentence.
Question: What is the name of india pakistan border?
Answer:
Reference Answer:
International Border (IB)
SQuAD_V2 Answer the following question concisely in one sentence.
Question: What was Warsaw’s population in 1901?
Answer:
Reference Answer:
711,988
SVAMP Answer the following question concisely in one sentence.
Context: A grocery store had 72 bottles of regular soda, 32 bottles of diet soda and 78 apples.
Question: How many more bottles than apple did they have? Answer:
Reference Answer:
26
TriviaQA Answer the following question concisely in one sentence.
Question: Traitor’s Gate is part of which building?
Answer:
Reference Answer:
tower of london
F.2 Datasets in long-form experiments
@@
We select DeepSeek-V3.1 and Gemini-3-Flash as representative SOTA LLMs at the time of writing. Due to the absence of fine-grained, long-form hallucination evaluation datasets for these two models, we construct four new datasets based on entities from FActScore Min et al. [2023] and PopQA Mallen et al. [2023].
@@
FActScore: This dataset is widely used for evaluating the factuality of biographies generated by LLMs, with entities sourced from Wikipedia. We select entities from the publicly released “unlabelled” portion and prompt each model with the query “Tell me a bio of wiki_entity:[subject]” to collect long-form generations.
@@
Annotation Process: For each model-dataset pair (four pairs in total), we employ a two-stage cross-validation annotation process. First, we generate factual claims for each sampled entity, following the processing pipeline detailed in Appendix B. Each claim that extracted from the greedily decoded response is then evaluated for factual accuracy according to the following criteria:
  • True: The claim is factually accurate and can be verified by the corresponding Wikipedia pages.
@@
To enhance annotation efficiency and consistency, we utilize two powerful reasoning models, GPT-5 and DeepSeek-R1, for automated labeling. A claim is automatically assigned a label only if both models produce identical classifications; otherwise, the claim is reviewed by human annotators for final decision. This hybrid annotation framework ensures high accuracy while enabling large-scale processing. In total, 7,407 generated claims have been annotated through this process.
We find that even such powerful LLMs exhibit significant hallucination rates, with 28% for DeepSeek-V3.1 and 25% for Gemini-3-Flash. Although hallucinations, like software vulnerabilities, cannot be completely eliminated at the current stage, effective detection mechanisms can mitigate associated risks and enhance the reliability of LLMs. Given that hallucinations may lead to serious real-world consequences, reliable uncertainty quantification is crucial for ensuring the responsible deployment of LLM-based systems, which highlights the important value of our research.
Below is the example of FactScore with DeepSeek-V3.1.
Dataset: FActScore, Model: DeepSeek-V3.1 Prompt:
Tell me a paragraph bio of wiki_entity: Danny Faure
  Greedily Decoded Generation:
Danny Faure is a Seychellois politician who served as the President of Seychelles from 2016 to 2020. He assumed office following the resignation of his predecessor, James Michel, and became the fourth president of the island nation. Faure, a member of the United Seychelles party, had previously held key ministerial positions, including Vice President and Minister of Finance. During his presidency, he focused on economic reforms, environmental conservation, and social development. After leaving office, he was appointed as the Special Envoy for the African Union on maritime security in the Indian Ocean. Faure is recognized for his contributions to Seychelles’ stability and sustainable development.
  Extracted Atomic Claims and Annotations: [T] Danny Faure is a Seychellois politician. [T] Danny Faure served as the President of Seychelles from 2016 to 2020. [T] Danny Faure assumed office following the resignation of his predecessor, James Michel. [T] Danny Faure became the fourth president of Seychelles. [T] Danny Faure is a member of the United Seychelles party. [T] Danny Faure previously held the position of Vice President of Seychelles. [F] Danny Faure previously held the position of Minister of Environment in Seychelles.
Annotation: Danny Faure previously held the position of Minister of Finance in Seychelles.
[F] During his presidency, Danny Faure focused on economic reforms.
Annotation: economic reforms were clearly recorded only during his tenure as finance minister in Wikipedia.
[T] During his presidency, Danny Faure focused on environmental conservation. [F] During his presidency, Danny Faure focused on social development.
Annotation: there is a lack of direct evidence to summarize the presidential term as "focusing on social development"
[F] Danny Faure was appointed as the Special Envoy for the African Union on maritime security in the Indian Ocean.
Annotation: there is no such position.
[T] Danny Faure is recognized for his contributions to Seychelles’ stability. [T] Danny Faure is recognized for his contributions to Seychelles’ sustainable development.
  Hallucination Analysis:
This serves as a typical example of how LLM long-form output interleaves true and false claims. While the model accurately describes most of the subject’s historical political career of Danny Faure, it hallucinates a nonexistent AU Special Envoy role and exaggerates attributions that are not supported by any authoritative references.
Appendix G Baselines Details
@@
G.1 Baselines in short-form experiments
@@
Length-normalized Predictive Entropy (LN-PE) Malinin and Gales [2021]. To calculate prediction entropy, we must obtain the probabilities that LLMs assign to generated token sequences. The probability of an entire sequence $s$ given context $x$ equals the product of probabilities for each token conditioned on previous tokens, with its log probability expressed as $\log P(s|x)=\sum_{i}\log P(s_{i}|s_{<i},x)$. $s_{i}$ is the $i$-th output token and $s_{<i}$ denotes all preceding tokens. Due to the conditional independence of token probabilities, longer sequences inherently have lower joint likelihood. The joint likelihood of a sequence decreases exponentially with length $L$, while its negative log probability increases linearly with $L$, causing longer sentences to contribute disproportionately to entropy. Length-normalized prediction entropy addresses this bias by normalizing the log probability by sequence length, using the arithmetic mean $\frac{1}{L}\sum_{i=1}^{L}\log P(s_{i}|s_{<i},x)$. This normalization effectively assumes that the uncertainty of the generated result is independent of sequence length.
P(True) Kadavath et al. [2022]. This method first prompts LLMs to generate multiple distinct answers, then presents this answer list alongside the greedy decoding response and a binary question: “Is this answer (a) correct or (b) incorrect?". The uncertainty score is determined by taking the negative of the probability that the LLMs answer "(a)" to this multiple-choice question. Following Farquhar et al. [2024], we enhance P(True) through few-shot prompting by incorporating ten randomly selected training examples formatted according to the described protocol and annotated with their true labels. This strategy represents a form of supervised in-context learning that leverages partial reference answer without necessitating model retraining.
SelfCheckGPT (SC) Manakul et al. [2023]. SelfCheckGPT (SC) is the first zero-resource hallucination detection solution that can be applied to black-box systems. SC-Prompt represents its highest-performing variant, which prompts the LLM to assess the semantic consistency between the target sentence and a set of randomly generated samples, thereby determining the presence of hallucinations. The specific procedure is as follows: given a sentence to be evaluated $r_{i}$ and $N$ randomly generated samples $S^{n}$ corresponding to the same query, a fixed prompt template is used to instruct the LLM to evaluate their semantic consistency. The LLM’s output is then converted into a numerical score $x$ in accordance with predefined rules. Finally, the inconsistency score for the sentence is obtained by averaging these individual scores. A score approaching 1.0 indicates a higher likelihood that the sentence contains hallucinations. The calculation formula is:
${SC}_{prompt}=\frac{1}{N}\sum_{n=1}^{N}x^{n},$ (21)
where $x_{i}^{n}$ is the mapping score corresponding to the $n$-th sample. In our experiment, we used the optimal variant ${SC}_{prompt}$ as the baseline.
Embedding Regression (ER) Farquhar et al. [2024]. Embedding Regression represents a typical supervised learning approach. Inspired by Kadavath et al. Kadavath et al. [2022], who developed predictors by fine-tuning proprietary language models on annotated QA datasets to assess whether target LLMs could correctly answer specific questions, [Farquhar et al., 2024] implemented a more efficient alternative. This approach directly extracts the final hidden layer states from LLMs and trains an Embedding Regression classifier to achieve equivalent predictive functionality without requiring model fine-tuning or ground-truth answers.
@@
Semantic Entropy (SE) Farquhar et al. [2024]. Semantic entropy aims to evaluate the uncertainty of LLMs regarding the meaning of their generated sequences. It first calculates the sum of probabilities for all token sequences that can be considered as expressing the same meaning. Given context $x$, for each semantic equivalence class $c$, its probability $P(c|x)$ is estimated through semantic clustering of generated sequences $s$:
@@
where $C_{i}$ are $M$ clusters extracted from the $N$ generations and $p^{\prime}(C_{i}\mid x)=\frac{p(C_{i}\mid x)}{\sum_{k}p(C_{k}\mid x)}$, which we refer to as $p(C_{i}\mid x)$ in the following for simplicity. DSE can be extended to cases where token likelihoods are not available by approximating $p(C_{i}\mid x)$ with the fraction of generated texts in each cluster, $p(C_{i}\mid x)\approx\sum_{i=1}^{N}\mathbb{I}(s_{i}\in C_{i})/N$.
Kernel Language Entropy (KLE)Nikitin et al. [2024]. To address the limitations of traditional semantic entropy, which only rely on semantic equivalence relations and fail to capture fine-grained semantic similarities, Nikitin et al. proposed Kernel Language Entropy (KLE). Its encodes the semantic similarities between texts generated by LLMs through a unit trace positive semidefinite kernel $K_{\text{sem}}$, and then use the von Neumann Entropy (VNE) to quantify the uncertainty of the semantic space represented by this kernel. For a unit trace positive definite matrix $A\in\mathbb{R}^{n\times n}$, its von Neumann Entropy, denoted as $\text{VNE}(A)$, is defined in the form of matrix trace operation. On this basis, the KLE is defined as the VNE of the semantic kernel:
@@
In our experiment, we used the optimal variant $\text{KLE}_{HEAT}$ as the baseline, which is a heat kernel over constructed semantic graph.
Semantic Graph Density (SGD)Li et al. [2025b]. [Li et al., 2025b] proposed Semantic Graph Density (SGD), which quantifies semantic consistency using graph density and adjusts edge contributions by incorporating answer probabilities. For $N$ white-box sampled answers $\{y^{(i)}\}_{i=1}^{N}$, pairwise semantic similarities $s_{ij}$ and length-normalized probabilities $P(y^{(i)}|x)$ are first computed to construct the semantic graph. SGD is defined as the negative value of semantic graph density, and the best-performing variant is version $\text{SGD}_{s+P}$, which optimizes edge contribution through a probability fusion strategy:
$\text{SGD}_{s+P}(x)=-\sum_{i<j}s_{ij}\cdot\mu(i,j),$ (27)
$\mu(i,j)=\theta\cdot\frac{1}{N(N-1)/2}+(1-\theta)\cdot\frac{P(y^{(i)}|x)P(y^{(j)}|x)}{\sum_{k<l}P(y^{(k)}|x)P(y^{(l)}|x)}.$ (28)
Where $\mu(i,j)$ is the edge contribution weight fusing prior uniform distribution and answer joint probability, and $\theta$ is a balancing hyperparameter. In our experiment, we used the optimal variant $\text{SGD}_{s+P}$ as the baseline.
G.2 Baselines in long-form experiments
Discrete Semantic Entropy (DSE) variant Farquhar et al. [2024]. For long-form generation, discrete semantic entropy operates through three key steps: (1) decomposing the text into atomic claims, (2) generating multiple possible questions that could trigger each claim in reverse, and (3) querying the original LLMs to produce new responses for each question, while including the original claim as a candidate. The final uncertainty estimate for each claim is derived by averaging the semantic entropy values across all associated questions. In our long-form experiments, we implement discrete semantic entropy according to the original paper’s best practices. Applying discrete semantic entropy to long-form generation introduces additional assumptions and complexities, and its computational cost increases with higher sampling. In contrast, SeSE achieves better performance with a more concise principle and lower cost.
Verbalized Uncertainty (VU) Mohri and Hashimoto [2024].
Verbalized Uncertainty prompts LLMs to directly express their confidence in a claim through natural language and maps confidence expressions in the LLMs’ output (e.g., “very confident", “100%", etc.) to numerical values. Uncertainty is quantified as the negative value of ther confidence score. Here, we mainly consider two variants:
  • Post-hoc Verbalized Uncertainty (PH-VU): This method elicits the verbalized confidence in a post-hoc manner after the entire claim set $C$ has been decomposed from generations. Specifically, we prompt an LLM to express its confidence about each claim $c\in C$ given multiple options such as “Unlikely (40%)”, “Even chance (50%)” etc. The specific prompt that we adapted from Mohri and Hashimoto [2024] is as following: